# The thermal view of $f(R)$ cosmology

### v1

@@ Abstract
A new thermal view of scalar-tensor gravity, in which general relativity is the zero-temperature state of gravity, is applied to the specific subclass of $f(R)$ gravity theories and, specifically, to spatially homogeneous and isotropic universes. Within the limits of application of the new thermal formalism, results on the convergence to Einstein cosmology (or lack thereof) are first obtained for general $f(R)$ theories, and then illustrated with power-law and Starobinsky $f(R)$ gravity.
@@ I Introduction
The simplest way of adding a new degree of freedom to GR is by making Newton’s constant $G$ a field varying across space and time, as proposed in the original Brans-Dicke theory [5] containing a gravitational scalar field $\phi$ and an effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$. Brans-Dicke theory was later generalized to “first generation” scalar tensor gravity [6, 7, 8, 9], still containing only a single extra scalar field $\phi$ in addition to GR, but making this scalar massive by endowing it with a potential $V(\phi)$, plus allowing the coefficient of the kinetic term of $\phi$ in the action to become a function of $\phi$ [6, 7, 8, 9].
@@
In the last decade, the quest for the most general scalar-tensor gravity with equations of motion of only second order has led to rediscovering Horndeski gravity [17], and to further generalize it with Degenerate Higher Order Scalar-Tensor (DHOST) theories. There is now a vast literature on scalar-tensor gravity and cosmology and it is difficult to gain a comprehensive view of all their aspects and observational constraints at different scales and regimes across the universe. A recent proposal [18, 19, 20, 21, 22, 23, 24, 25, 26] views scalar-tensor gravity as a thermally excited state of GR, which corresponds to the “zero-temperature” state of gravity. The approach of scalar-tensor gravity to GR occurring, for example, in the early universe [27, 28] (with its complications [29]), is seen as the analogue of heat conduction in an effective dissipative fluid associated with the scalar degree of freedom $\phi$. A new formalism describing these situations (under certain restrictions) is currently under development and is being tested on various scalar-tensor theories and their exact solutions. Here we continue this work, applying this thermal picture (described below) to $f(R)$ cosmology. In this context, the new thermal view of scalar-tensor gravity unifies several theoretical results that appeared in a rather fragmented literature over the years and predicts when, and under what conditions, $f(R)$ universes converge to, or depart from, GR ones. Two basic ideas of the thermal view of scalar-tensor gravity are: i) gravity is “hot” (i.e., deviates from GR) in strong gravity, in particular near spacetime singularities; ii) the expansion of 3-space “cools” gravity, making it approach GR analogously to heat conduction in a dissipative fluid. Although these basic ideas prove true in simplified situations, more involved regimes involve the presence of “heat sources” or “sinks” in the relevant equations, which complicate the basic picture. These more involved situations occur in $f(R)$ cosmology, as discussed in the following sections.
@@
$u^{a}=\frac{\nabla^{a}\phi}{\sqrt{-\nabla_{c}\phi\nabla^{c}\phi}}$ (1.6)
@@
where $\tau$ is the proper time along the effective $\phi$-fluid lines.
@@
The recent Ref. [25] advances the thermal picture with the study of scalar-tensor situations in which $\Box\phi=0$, in which case the $\left(\Theta,{\cal K}{\cal T}\right)$ plane (not a phase space) provides a convenient representation of the behaviour of gravity. If the expansion scalar $\Theta$ is negative, or if the system begins at a point above the critical half-line $8\pi{\cal K}{\cal T}=\Theta>0$, then gravity can only “heat up” and diverge away from GR. If instead $\Theta>0$ and the system begins at a point below this critical half-line (i.e., with $8\pi{\cal K}{\cal T}<\Theta$), gravity can only “cool” and converge to GR [25]. This picture survives if conformal matter with $T^{(m)}=0$, but no potential, is added to the picture. Here we study vacuum $f(R)$ gravity, which unavoidably adds a potential for $\phi=f^{\prime}(R)$.
Section II reviews $f(R)$ gravity and cosmology, while the following sections apply the thermal view to it.
@@ III Thermal view of FLRW cosmology in general $f(R)$ gravity
First, it is crucial that the dynamics near a Big Bang singularity is dominated by gravity instead of matter. This behaviour is quite plausible [45, 26], but is not a given and should be investigated case by case. When it turns out to be true, the dynamics reduces to that of vacuum $f(R)$ gravity.
We now know that, for vacuum or conformal matter, whether the conjecture of [41] is satisfied or not depends on the initial conditions. Expanding FLRW universes which start out with $8\pi{\cal K}{\cal T}>\Theta>0$, plus contracting ones with $\Theta<0$, inescapably depart from GR, but solutions that start out with $8\pi{\cal K}{\cal T}<\Theta$ do approach GR [25]. Solutions with $8\pi{\cal K}{\cal T}=\Theta$ are fixed points with constant ${\cal K}{\cal T}$ and constant $\Theta$, but they are thermally unstable [25]. The thermal picture settles the conjecture of [41] for vacuum or conformal matter with a completely different approach. For more general forms of matter, for which the term $\frac{T^{\mathrm{(m)}}}{(2\omega+3)\phi}$ in Eq. (LABEL:evolution_general2) is non-vanishing, the thermal picture is more complicated.
In their perturbative analysis, the authors of [46, 47, 41] find $k=0$ universes that are regular at $t=0$ (contrary to GR universes that have a Big Bang singularity) and converge to GR solutions as $t\to+\infty$, apparently contradicting the general thermal view of scalar-tensor gravity. The latter states that universes starting out with $8\pi{\cal K}{\cal T}>\Theta>0$ (which they do if they initially depart drastically from GR) always run away from GR in the future and cannot approach it [25]. However, these universes of [46, 47, 41] are obtained for negative values of the parameter $\lambda^{2}\equiv\frac{f_{0}^{\prime}}{3f_{0}^{\prime\prime}}$ (the zero subscript denotes quantities evaluated on the unperturbed solution). Unbeknown to the authors of Refs. [46, 47, 41] working in the 1970s and 1990s, we now know that in $f(R)$ gravity it must be $f^{\prime}(R)>0$ and $f^{\prime\prime}(R)>0$ [14, 15, 16], therefore, the region $\lambda^{2}<0$ is physically irrelevant. [^2] The same conclusion applies to the solutions (13), (15), (19), (21), and (27) of Ref. [41].
@@
  • pure $R^{2}$ gravity (which is in some sense pathological [51, 52, 53, 54]), is the threshold between these two behaviours, corresponding to the quadratic potential $V(\phi)=\alpha\phi^{2}/2$, which disappears from the field equation (1.4) for $\phi$. Then the electrovacuum evolution equation for ${\cal K}{\cal T}$ reduces to
@@ IV FLRW universes in polynomial and power-law $f(R)$ gravity
Figure 1: In the region of the $\left(\Theta,{\cal K}{\cal T}\right)$ plane between the half-line ${\cal K}{\cal T}=\Theta/(8\pi)$ (black, solid upper line) and the positive $\Theta$-axis, gravity “cools” and approaches GR. Outside of this region, gravity “heats up” and departs from GR. The red, dashed line corresponds to the solution (5.5)-(5.6) of vacuum $R^{n}$ gravity; the blue, dash-dotted line corresponds to the solution (5.15)-(5.17) of $R^{n}$ gravity with radiation, In both cases, $n=1.3$ and gravity moves towards the origin and GR.
@@ V Power-law $f(R)$
The left-hand side is $d\left({\cal K}{\cal T}\right)/dt=-\frac{(n-1)}{4\pi t^{2}}$, while the right hand-side results from the competition of cooling and heating terms:
@@
which, of course, equals the left-hand side and shows how the heating term
$I\equiv{\cal K}{\cal T}\left(8\pi{\cal K}{\cal T}-\Theta\right)=\frac{(n-1)(n-4)}{8\pi t^{2}}$ (5.22)
(for $1<n<4$) is smaller than the cooling term
$II\equiv\frac{(2-n)(n-1)}{8\pi t^{2}}\,.$ (5.23)
@@
$\left|\frac{I}{II}\right|=\left|\frac{n-4}{n-2}\right|$ (5.24)
is never larger than unity so the “cooling” term prevails for $1<n<4$.
@@ V.1 Pure $R^{2}$-gravity
If $R_{\infty}=0$, we have an asymptotic Minkowski space corresponding to $\phi=2R\simeq 0$ and infinite effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$, which is clearly away from GR. Indeed, it is well known that $R^{2}$ gravity does not admit a Newtonian limit around Minkowski space [51] (although it does admit one around de Sitter space [110]). Similarly, a Hamiltonian analysis reveals the absence of propagating degrees of freedom around Minkowski space, or any background with $R=0$, in the linear approximation (although the full theory propagates three degrees of freedom) [112].
@@ VI.1 Limit $R\to 0$
and GR is approached as $R(t)$ approaches its horizontal asymptote $R=0$. At this stage, if $R\to 0$ and assuming that it stays small, Ref. [25] and Fig. 1 tell us that initial conditions such that the dynamics begins below the half-line $8\pi{\cal K}{\cal T}=\Theta>0$ guarantee that gravity converges to GR while, if the universe begins above this line, or contracts with $\Theta<0$, it should always run away from GR. So why does Starobinsky gravity always converge to GR as the universe expands? The key is that, as $R\to 0$, Eq. (6.10) yields
@@ VI.2 Strong gravity regime $R\to\infty$
In quadratic gravity there are spatially flat, radiation-dominated ($T^{\mathrm{(m)}}=0$) FLRW universes that avoid the initial Big Bang singularity and diverge in the future, $R(t)\to\infty$ as $t\to+\infty$ [46, 116, 117, 40]. Clearly, as $R\to\infty$ it is $f(R)=R+\alpha R^{2}\simeq\alpha R^{2}$ and Starobinsky gravity departs drastically from GR (but the perturbative approach leading to Starobinsky gravity will break down at some point in this limit). These solutions are found to be unstable and, therefore, unphysical [40]. In any case, they do not begin with $R=0$, but with a finite $R$. For these solutions, if $R(0)$ is larger than a critical value $R_{*}$, the solutions diverge as $t\to+\infty$, that is they diverge from GR. If they begin with $R(0)<R_{*}$, then $R(t)\to 0^{+}$ as $t\to+\infty$, approaching GR. The thermal view can be applied to these solutions with decreasing $R(t)$ and their behaviour matches the results of the thermal view of scalar-tensor gravity [25] (applied to the specific subclass of Starobinsky gravity) and could have been predicted had this formalism been available in 1971. The thermal formalism cannot be applied to the solutions with $\dot{R}>0$.
@@
where matter is a radiation fluid, restricting themselves to $k=0$ FLRW solutions. Therefore, Eq. (6.4) applies and, on FLRW solutions, the theory can be recast as Starobinsky gravity $f(R)=R+\bar{\alpha}R^{2}$, where $\bar{\alpha}=\eta\left(1+\alpha/3\right)$. Nariai and Tomita looked for bouncing universes that avoid the Big Bang singularity [116]. Clearly, this is non-GR behaviour because the radiation fluid always obeys the energy conditions and, in GR, a bounce with $\dot{H}>0$ requires the violation of the energy condition since $\dot{H}=-4\pi\left(\rho+P\right)$. This physics can be provided by the effective $\phi$-fluid which, being a fictitious fluid, can violate all of the energy conditions applicable to a real fluid. Nariai and Tomita integrated numerically the field equation for the scale factor $a(t)$ and found two types of solutions, which both begin by contracting ($\Theta=3H<0$) and reach an instant of time (which we will call $t=0$) at which $\dot{a}=0$ and $\Theta=3\dot{a}/a=0$ [116]. At times $t>0$, one class of solutions expands forever, asymptoting to the GR solution $a(t)\simeq\sqrt{t}$ as $t\to+\infty$, which means that ${\cal K}{\cal T}\to 0^{+}$ while $R\to 0$. The second class of solutions contracts again at $t>0$, ending in a Big Crunch with $\Theta\to$-$\infty$ and $R\to\infty$ at a finite time $t_{0}>0$ [116]. This means that $\ddot{a}$ changes from positive during the first contraction, to zero at $t=0$, to negative for $0<t<t_{0}$. This second type of solutions clearly departs from GR, but the thermal formalism is not applicable. In fact, for large $|R|$ it is $\phi\simeq 2\alpha R$, which implies that it must be $R>0$. At the finite time singularity $R$ diverges so it must be $R\to+\infty$. However, the Nariai-Tomita second class of solutions has $\dot{a}=0$ and $\ddot{a}=0$ at $t=0$, hence $R$ increases for times $0<t<t_{0}$, hence $\dot{R}<0$ is not satisfied, $\dot{\phi}>0$, and the thermal formalism is not applicable.

### v2

@@ Abstract
A new thermal view of scalar-tensor gravity, in which general relativity is the zero-temperature state of gravity, is applied to the subclass of $f(R)$ gravity theories and, specifically, to spatially homogeneous and isotropic universes. Within the limits of application of the new thermal formalism, results on the convergence to Einstein cosmology (or lack thereof) are first obtained for general $f(R)$ theories, and then illustrated with power-law and Starobinsky $f(R)$ gravity.
@@ I Introduction
The simplest way of adding a new degree of freedom to GR is by making Newton’s constant $G$ a field varying across space and time, as proposed in the original Brans-Dicke theory [5] containing a gravitational scalar field $\phi$ and an effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$. Brans-Dicke theory was later generalized to “first generation” scalar-tensor gravity [6, 7, 8, 9], still containing only a single extra scalar field $\phi$ in addition to GR, but making this scalar massive by endowing it with a potential $V(\phi)$, plus allowing the coefficient of the kinetic term of $\phi$ in the action to become a function of $\phi$ [6, 7, 8, 9].
@@
In the last decade, the quest for the most general scalar-tensor gravity with equations of motion of only second order has led to rediscovering Horndeski gravity [17], and to further generalize it with Degenerate Higher Order Scalar-Tensor (DHOST) theories. There is now a vast literature on scalar-tensor gravity and cosmology and it is difficult to gain a comprehensive view of all their aspects and observational constraints at different scales and regimes across the universe. A recent proposal ([18, 19, 20, 21, 22, 23, 24, 25, 26], see also [27]) views scalar-tensor gravity as a thermally excited state of GR, which corresponds to the “zero-temperature” state of gravity. The approach of scalar-tensor gravity to GR occurring, for example, in the early universe [28, 29] (with its complications [30]), is seen as the analogue of heat conduction in an effective dissipative fluid associated with the scalar degree of freedom $\phi$. A new formalism describing these situations (under certain restrictions) is currently under development and is being tested on various scalar-tensor theories and their exact solutions. Here we continue this work, applying this thermal picture (described below) to $f(R)$ cosmology. In this context, the new thermal view of scalar-tensor gravity unifies several theoretical results that appeared in a rather fragmented literature over the years and predicts when, and under what conditions, $f(R)$ universes converge to, or depart from, GR ones. Two basic ideas of the thermal view of scalar-tensor gravity are: i) gravity is “hot” (i.e., deviates from GR) in strong gravity, in particular near spacetime singularities; ii) the expansion of 3-space “cools” gravity, making it approach GR analogously to heat conduction in a dissipative fluid. Although these basic ideas prove true in simplified situations, more involved regimes involve the presence of “heat sources” or “sinks” in the relevant equations, which complicate the basic picture. These more involved situations occur in $f(R)$ cosmology, as discussed in the following sections.
The new thermal formalism basically consists of an analogy between the dynamics of gravity and heat dissipation in a relativistic imperfect fluid. This analogy has limitations in its regime of applicability and, like all analogies, when using it one should not confuse its two sides, which describe physically different systems. However, since there are two sides to the analogy, one can use one of them (the most intuitive and better known one) to understand the other (less known) side, which is what we do here.
@@
$u^{a}\equiv\frac{\nabla^{a}\phi}{\sqrt{-\nabla_{c}\phi\nabla^{c}\phi}}$ (1.6)
@@
where $\tau$ is the proper time along the effective $\phi$-fluid lines and $\Theta\equiv\nabla_{c}u^{c}$ is its expansion scalar.
@@
The recent Ref. [25] advances the thermal picture with the study of scalar-tensor situations in which $\Box\phi=0$, in which case the $\left(\Theta,{\cal K}{\cal T}\right)$ plane (not a phase space) provides a convenient representation of the behavior of gravity. If the expansion scalar $\Theta$ is negative, or if the system begins at a point above the critical half-line $8\pi{\cal K}{\cal T}=\Theta>0$, gravity either “heats up” diverging away from GR, or else this line is crossed with horizontal tangent, after which gravity cools forever [25, 36].
If instead $\Theta>0$ and the system begins at a point below this critical half-line (i.e., with $8\pi{\cal K}{\cal T}<\Theta$), gravity can only “cool” and converge to GR [25]. This picture survives if conformal matter with $T^{(m)}=0$, but no potential, is added to the picture. Here we study vacuum $f(R)$ gravity, which unavoidably adds a potential for $\phi=f^{\prime}(R)$.
Section II reviews $f(R)$ gravity and cosmology, while the following sections apply the thermal view to it. Several exact solutions of $f(R)$ gravity are known (e.g., [37, 38, 39, 40, 41]) and we use some of them as examples.
@@ III Thermal view of FLRW cosmology in general $f(R)$ gravity
First, it is crucial that the dynamics near a Big Bang singularity is dominated by gravity instead of matter. This behavior is quite plausible [52, 26], but is not a given and should be investigated case by case. When it turns out to be true, the dynamics reduces to that of vacuum $f(R)$ gravity.
We now know that, for vacuum or conformal matter, whether the conjecture of [48] is satisfied or not depends on the initial conditions. Expanding FLRW universes which start out with $8\pi{\cal K}{\cal T}>\Theta>0$, plus contracting ones with $\Theta<0$, depart from GR, but solutions that start out with $8\pi{\cal K}{\cal T}<\Theta$ do approach GR [25]. Solutions with $8\pi{\cal K}{\cal T}=\Theta$ and with constant ${\cal K}{\cal T}$ and constant $\Theta$ are fixed points, but they are thermally unstable [25]. The thermal picture settles the conjecture of [48] for vacuum or conformal matter with a completely different approach. For more general forms of matter, for which the term $\frac{T^{\mathrm{(m)}}}{(2\omega+3)\phi}$ in Eq. (LABEL:evolution_general2) is non-vanishing, the thermal picture is more complicated.
In their perturbative analysis, the authors of [53, 54, 48] find $k=0$ universes that are regular at $t=0$ (contrary to GR universes that have a Big Bang singularity) and converge to GR solutions as $t\to+\infty$, apparently contradicting the general thermal view of scalar-tensor gravity. The latter states that universes starting out with $8\pi{\cal K}{\cal T}>\Theta>0$ (which they do if they initially depart drastically from GR) run away from GR as long as $8\pi{\cal K}{\cal T}$ remains larger than $\Theta$ [25]. However, these universes of [53, 54, 48] are obtained for negative values of the parameter $\lambda^{2}\equiv\frac{f_{0}^{\prime}}{3f_{0}^{\prime\prime}}$ (the zero subscript denotes quantities evaluated on the unperturbed solution). Unbeknownst to the authors of Refs. [53, 54, 48] working in the 1970s and 1990s, we now know that in $f(R)$ gravity it must be $f^{\prime}(R)>0$ and $f^{\prime\prime}(R)>0$ [14, 15, 16], therefore, the region $\lambda^{2}<0$ is physically irrelevant. [^2] The same conclusion applies to the solutions (13), (15), (19), (21), and (27) of Ref. [48].
@@
  • pure $R^{2}$ gravity (which is in some sense pathological [58, 59, 60, 61]), is the threshold between these two behaviors, corresponding to the quadratic potential $V(\phi)=\alpha\phi^{2}/2$, which disappears from the field equation (1.4) for $\phi$. Then the electrovacuum evolution equation for ${\cal K}{\cal T}$ reduces to
@@ IV FLRW universes in polynomial and power-law $f(R)$ gravity
Figure 1: In the region of the $\left(\Theta,{\cal K}{\cal T}\right)$ plane between the half-line ${\cal K}{\cal T}=\Theta/(8\pi)$ (black, solid upper line) and the positive $\Theta$-axis, gravity “cools” and approaches GR. Outside of this region, gravity “heats up” and departs from GR. The red, dashed line corresponds to the solution (5.5)-(5.6) of vacuum $R^{n}$ gravity; the blue, dash-dotted line corresponds to the solution (5.15)-(5.17) of $R^{n}$ gravity with radiation. In both cases, $n=1.3$ and gravity moves towards the origin and GR.
@@ V Power-law $f(R)$
The left-hand side is $d\left({\cal K}{\cal T}\right)/dt=-\frac{(n-1)}{4\pi t^{2}}$, while the right- hand side results from the competition of cooling and heating terms:
@@
which, of course, equals the left-hand side and shows how the heating term (e.g., for $1<n<2$)
$I\equiv 8\pi\left({\cal K}{\cal T}\right)^{2}+\frac{(2-n)(n-1)}{8\pi t^{2}}=\frac{(n-1)(3n-2)}{8\pi t^{2}}$ (5.22)
is smaller than the cooling term
$II\equiv\frac{-3n(n-1)}{8\pi t^{2}}\,.$ (5.23)
@@
$\left|\frac{I}{II}\right|=\left|\frac{3n-2}{3n}\right|$ (5.24)
is never larger than unity so the “cooling” term prevails for $1<n<2$.
@@ V.1 Pure $R^{2}$-gravity
If $R_{\infty}=0$, we have an asymptotic Minkowski space corresponding to $\phi=2R\simeq 0$ and infinite effective gravitational coupling $G_{\mathrm{eff}}\simeq 1/\phi$, which is clearly away from GR. Indeed, it is well known that $R^{2}$ gravity does not admit a Newtonian limit around Minkowski space [58] (although it does admit one around de Sitter space [117]). Similarly, a Hamiltonian analysis reveals the absence of propagating degrees of freedom around Minkowski space, or any background with $R=0$, in the linear approximation (although the full theory propagates three degrees of freedom) [118]. Counting the propagating degrees of freedom is far from trivial and depends on the background [119, 120, 121].
@@ VI.1 Limit $R\to 0$
and GR is approached as $R(t)$ approaches its horizontal asymptote $R=0$. At this stage, if $R\to 0$ and assuming that it stays small, Ref. [25] and Fig. 1 tell us that initial conditions such that the dynamics begins below the half-line $8\pi{\cal K}{\cal T}=\Theta>0$ guarantee that gravity converges to GR while, if the universe begins above this line, or contracts with $\Theta<0$, it should run away from GR. So why does Starobinsky gravity always converge to GR as the universe expands? The key is that, as $R\to 0$, Eq. (6.10) yields
@@ VI.2 Strong gravity regime $R\to\infty$
In quadratic gravity there are spatially flat, radiation-dominated ($T^{\mathrm{(m)}}=0$) FLRW universes that avoid the initial Big Bang singularity and diverge in the future, $R(t)\to\infty$ as $t\to+\infty$ [53, 126, 127, 47]. Clearly, as $R\to\infty$ it is $f(R)=R+\alpha R^{2}\simeq\alpha R^{2}$ and Starobinsky gravity departs drastically from GR (but the perturbative approach leading to Starobinsky gravity will break down at some point in this limit). These solutions are found to be unstable and, therefore, unphysical [47]. In any case, they do not begin with $R=0$, but with a finite $R$. For these solutions, if $R(0)$ is larger than a critical value $R_{*}$, the solutions diverge as $t\to+\infty$, that is they diverge from GR. If they begin with $R(0)<R_{*}$, then $R(t)\to 0^{+}$ as $t\to+\infty$, approaching GR. The thermal view can be applied to these solutions with decreasing $R(t)$ and their behavior matches the results of the thermal view of scalar-tensor gravity [25] (applied to the specific subclass of Starobinsky gravity) and could have been predicted had this formalism been available in 1971. The thermal formalism cannot be applied to the solutions with $\dot{R}>0$.
@@
where matter is a radiation fluid, restricting themselves to $k=0$ FLRW solutions. Therefore, Eq. (6.4) applies and, on FLRW solutions, the theory can be recast as Starobinsky gravity $f(R)=R+\bar{\alpha}R^{2}$, where $\bar{\alpha}=\eta\left(1+\alpha/3\right)$. Nariai and Tomita looked for bouncing universes that avoid the Big Bang singularity [126]. Clearly, this is non-GR behavior because the radiation fluid always obeys the energy conditions and, in GR, a bounce with $\dot{H}>0$ requires the violation of the energy condition since $\dot{H}=-4\pi\left(\rho+P\right)$. This physics can be provided by the effective $\phi$-fluid which, being a fictitious fluid, can violate all of the energy conditions applicable to a real fluid. Nariai and Tomita integrated numerically the field equation for the scale factor $a(t)$ and found two types of solutions, which both begin by contracting ($\Theta=3H<0$) and reach an instant of time (which we will call $t=0$) at which $\dot{a}=0$ and $\Theta=3\dot{a}/a=0$ [126]. At times $t>0$, one class of solutions expands forever, asymptoting to the GR solution $a(t)\simeq\sqrt{t}$ as $t\to+\infty$, which means that ${\cal K}{\cal T}\to 0^{+}$ while $R\to 0$. The second class of solutions contracts again at $t>0$, ending in a Big Crunch with $\Theta\to$-$\infty$ and $R\to\infty$ at a finite time $t_{0}>0$ [126]. This means that $\ddot{a}$ changes from positive during the first contraction, to zero at $t=0$, to negative for $0<t<t_{0}$. This second type of solutions clearly departs from GR, but the thermal formalism is not applicable. In fact, for large $|R|$ it is $\phi\simeq 2\alpha R$, which implies that it must be $R>0$. At the finite time singularity $R$ diverges so it must be $R\to+\infty$. However, the Nariai-Tomita second class of solutions has $\dot{a}=0$ and $\ddot{a}=0$ at $t=0$, hence $R$ increases for times $0<t<t_{0}$, hence $\dot{R}<0$ is not satisfied, $\dot{\phi}>0$, and the thermal formalism is not applicable.